The distance from earth surface to moon surface is 3.8x105 km. Calculate areal spread and angular spread for a laser of wavelength 632.8 nm and aperture 0.5 mm when it reaches moon.


1
Expert's answer
2021-05-03T10:29:00-0400

Angular Spread, "\theta=\dfrac{wavelength(\lambda)}{Aperture(a)}=\dfrac{632.8\times10^{-9}}{0.5\times10^{-3}}"

"\theta=1265.6\times10^{-6}"


Axial spread "=d\times\theta=(3.8\times10^5)\times(1265.6\times10^{-6})=480.928"